TEMPERATURE DEPENDENCES OF MECHANISMS RESPONSIBLE FOR THE WATER VAPOR CON-TINUUM

\underline{Q. MA}, NASA/Goddard Institute for Space Studies and Department of Applied Physics and Applied methamatics, Columbia University, 2880 Broadway, New York, NY 10025; R. H. TIPPING, Department of Physics and Astronomy, University of Alabama, Tuscaloosa, AL 35487; C. LEFOR-ESTIER, Institute Charles Gerhardt CNRS-5253, CC1501, Université Montpellier II, 34095 Montpellier, France.

The water-vapor continuum plays an important role in the radiation balance in the Earth’s atmosphere. While this absorption has been known for a long time, the physical mechanism responsible is still an open problem. We have recently calculated theoretically both the magnitude and temperature dependence for the three mechanism that have been suggested: the far-wings of allowed lines, collision-induced absorption, and water dimers^{a}^{,}^{b}. All three mechanisms depend quadratically on the number density of H_{2}O for the self-continuum, and on the product of the densities for the foreign-continuum. However, these three mechanisms have quite differences on the temperature as we will discuss. This analysis may provide us with a method to assess their relative importance in ambient atmospheric measurements.

aQ. Ma, R. H. Tipping, and C. Leforestier, J. Chem. Phys. 128, 124313 (2008).

bC. Leforestier, R. H. Tipping, and Q. Ma, to be submitted for publication.